# spade: Synthesizing Assertions for Large Language Model Pipelines

### v1

@@ 4.3. Results and Discussion
Subsumption as a Means for Reducing Redundancy. Several pipelines exhibit a large discrepancy between functions selected in $\textsc{spade}_{\_}{\text{base}}$ and $\textsc{spade}_{\_}{\text{sub}}$, which occurs when there are many redundant candidates. For example, in the codereviews pipeline’s 8 prompt versions, the developer iterated several times on the instruction to give a clear and concise review, resulting in five assertions that check the same thing (two of which are shown below): Since all the five assertions meet the FFR constraint, individually, $\textsc{spade}_{\_}{\text{base}}$ would select them all, which is undesirable because they all do the same thing, but $\textsc{spade}_{\_}{\text{sub}}$ would select the one most compatible with the FFR constraint, as long as subsumption is assessed correctly. On the flip side, while assessing subsumption, the LLM may not recall all subsumptions, so $\textsc{spade}_{\_}{\text{sub}}$ may have duplicate assertions. For example, the codereviews pipeline contains assertions titled assert_includes_code_improvementv1 and assert_includes_code_improvementv2. $\boldsymbol{\alpha}$ and $\boldsymbol{\tau}$ Threshold Sensitivity. The feasibility of solutions from the ILP solver in spade is dependent on the chosen $\alpha$ and $\tau$ thresholds. If a feasible solution is not found, developers may need to adjust these values in a binary search fashion. In our case, all 9 LLM pipelines yielded feasible solutions with $\alpha=0.6$ and $\tau=0.25$. However, the small size of $E^{\prime}$ makes $\textsc{spade}_{\_}{\text{cov}}$ particularly sensitive to $\alpha$. In the pipelines, we observed that between one and five assertions covered 60% of $E^{\prime}$’s failures. For example, $\textsc{spade}_{\_}{\text{cov}}$ selected only one assertion for the emails pipeline: If $E^{\prime}$ is exhaustive of failure modes and representative of the distribution of failures (e.g., for the emails pipeline, most failures are actually due to the response lacking an encouragement to contact the company), $\textsc{spade}_{\_}{\text{cov}}$ might be a satisfactory solution. However, our $E^{\prime}$ datasets clearly were not exhaustive, considering that $\textsc{spade}_{\_}{\text{sub}}$ always chose additional assertions. $\textsc{spade}_{\_}{\text{sub}}$ is less sensitive to $\alpha$, as it explicitly selects assertions based on their potential to cover new failures (i.e., subsumption) without exceeding the FFR, even if the constraint on coverage is no longer tight. FFR Tradeoffs. Considering that the difference between the fraction of functions selected for $\textsc{spade}_{\_}{\text{base}}$ and $\textsc{spade}_{\_}{\text{sub}}$ is less than 10% for three LLM pipelines, one may wonder if the complexity of $\textsc{spade}_{\_}{\text{sub}}$ is worth it. $\textsc{spade}_{\_}{\text{sub}}$ is generally preferable because $\textsc{spade}_{\_}{\text{base}}$ fails to consistently meet the FFR threshold $\tau$. We observed that as prompt versions increase, so do the number of assertions, impacting $\textsc{spade}_{\_}{\text{base}}$ adversely. The worst-case FFR of a set is the sum of individual FFRs, as shown in Equation 2. Hence, with a large number of independent assertions, the total FFR is likely to surpass the threshold. This issue is evident in the fashion and lecturesummaries pipelines, where despite each of the 67 and 32 assertions meeting FFR constraints individually, the total FFR for $\textsc{spade}_{\_}{\text{base}}$ reaches 0.88 and 0.53, respectively. In practice, if spade were to be deployed in an interactive system, where spade could observe each LLM call in real-time (e.g., as a wrapper around the OpenAI API), the multitude of prompt versions further necessitates filtering assertions based on overall FFR. This underscores the need for the more complex $\textsc{spade}_{\_}{\text{cov}}$ or $\textsc{spade}_{\_}{\text{sub}}$ approaches.
@@ 5. Related Work
Prompt Engineering. For both nontechnical (Zamfirescu-Pereira et al., 2023) and technical users (Si et al., 2022; Parameswaran et al., 2023), prompt engineering is hard for several reasons: small changes in prompt phrasing (Liu et al., 2023; Arora et al., 2022) or the order of instructions or contexts (Lu et al., 2021) can significantly affect outputs. Moreover, as LLMs change under the hood of the API (i.e., prompt drift), outputs can change without developer awareness (Chen et al., 2023). Tools and papers are emerging to aid in prompt management and experimentation, and are even using LLMs to write prompts (Arawjo et al., 2023; Wu et al., 2022b, a; Cheng et al., 2023; Zhou et al., 2022). Moreover, deployed prompts introduce new challenges, like “balancing more context with fewer tokens” and “wrangling prompt output” to meet user-defined criteria (Parnin et al., 2023). Our work doesn’t focus explicitly on helping developers create better prompts, but it could indirectly support developers in improving prompts through recommended assertions.
ML and LLM Evaluation. Evaluating and monitoring deployed ML models is known to be challenging (Paleyes et al., 2022; Shankar et al., 2022b). Evaluating LLMs in a deployed setting is even more challenging because LLMs are typically used for generative tasks, where the outputs are free-form (Dou et al., 2021). Some LLM pipeline types, like question-answering with retrieval-augmented generation pipelines (Lewis et al., 2020), can use standardized, automated metrics (Es et al., 2023; Saad-Falcon et al., 2023), but others face challenges due to unknown metrics and lack of labeled datasets (Xiao et al., 2023; Chang et al., 2023; Parnin et al., 2023). Typically, organizations rely on human evaluators for LLM outputs (Gehrmann et al., 2023; Wang et al., 2023b; Parnin et al., 2023), but recent studies suggest LLMs can self-evaluate effectively with detailed “scorecards” (Chan et al., 2023; Zhang et al., 2023; Kim et al., 2023). However, writing these scorecards can be challenging (Parnin et al., 2023), motivating auto-generated evaluators.
@@
Testing in ML Pipelines. ML pipelines are hard to manage in production. Much of the literature on ML testing is geared towards validating structured data, through analyzing data quality (Hynes et al., 2017; Breck et al., 2019; Schelter et al., 2018; Shankar et al., 2023a) or provenance (Schelter et al., 2023; Namaki et al., 2020). Platforms for ML testing typically offer automated experiment tracking and prevention against overfitting (Renggli et al., 2019; Agrawal et al., 2019), as well as tools for data distribution debugging (Grafberger et al., 2022). Model-specific assertions typically require human specification (Kang et al., 2020), or at least large amounts of data to train learned assertions (Kang et al., 2022). LLM chains or pipelines are a new class of ML pipelines, and LLMs themselves can generate assertions with little data. A recent study highlights the difficulty of testing LLM pipelines for “copilot”-like products: developers want to ensure accuracy while avoiding excessive resource use, such as running hundreds of assertions (Parnin et al., 2023)—motivating our assertion filtering approach.
6. Conclusion
Our work introduced a new problem of auto-generating assertions to catch failures in LLM pipelines, as well as a solution that comprises of two components: first, it synthesizes candidate assertions; then, it filters them. We proposed a taxonomy of prompt edits for assertion synthesis, demonstrating its potential via integration and deployment with LangChain. We expressed the selection of an optimal set of assertions for covering failures with high accuracy as an Integer Linear Program (ILP). We proposed assertion subsumption to cover failures in data-scarce scenarios and incorporated this into our ILP. Our auto-generating assertion system, spade, was evaluated on nine real-world data-generating LLM pipelines. We have made our code and datasets public for further research and analysis.
@@ Appendix A Notation
$\mathcal{P}$
A prompt template (i.e., a string with placeholders, intended to be submitted to an LLM)
$\mathcal{P}_{\_}i$
The $i$th version of a prompt template
$\Delta\mathcal{P}_{\_}i$
The diff or changes between consecutive versions of a prompt template
$E$
The set of all hypothetical example runs for an LLM pipeline
$E^{\prime}$
A small set of example runs for an LLM pipeline (with labels denoting whether responses were good) that we can observe, where $E^{\prime}\subset E$
$n$
The number of examples in $E^{\prime}$
$e_{\_}i$
The $i$th example in $E^{\prime}$
$f$
An arbitrary assertion function
$F$
The set of all candidate assertions from Section 2
$m$
The number of assertions in $F$
$F^{\prime}$
A subset of $F$ selected as a minimal set of assertions
$G$
Set of functions in $F\setminus F^{\prime}$ not subsumed by $F^{\prime}$ (as defined by Definition 3.3)
$\hat{y}_{\_}i$
Binary indicator of whether example $e_{\_}i$ satisfies all assertions in $F^{\prime}$
$y_{\_}i$
Binary indicator of whether a developer considers the LLM pipeline response for example $e_{\_}i$ good (i.e., success) or bad (i.e., failure)
$\alpha$
Threshold for coverage of failing examples; $F^{\prime}$ should cover at least $\alpha$ of failing examples
$\tau$
Threshold for False Failure Rate (FFR); $F^{\prime}$ should fail no more than $\tau$ good examples
$M$
$n\times m$ matrix representing the results of assertions on examples in $E^{\prime}$
$K$
$m\times m$ matrix representing subsumption relationships between functions
$x_{\_}j$
ILP variable indicating inclusion of $f_{\_}j$ in $F^{\prime}$
$w_{\_}{ij}$
ILP variable representing if $F^{\prime}$ denotes $e_{\_}i$ as a failure
$u_{\_}i$
ILP variable indicating if a failed example is covered
$z_{\_}i$
ILP variable defining if $F^{\prime}$ incorrectly marks a successful example as a failure (false failure)
$r_{\_}j$
ILP variable indicating if a function is subsumed by any function in $F^{\prime}$
$s_{\_}j$
ILP variable representing functions in $F\setminus F^{\prime}$ not subsumed by $F^{\prime}$
@@ Appendix B LangChain Prompt Hub and Taxonomy of Prompt Deltas
Let concepts be the parsed categories identified by the previous prompt. The prompt for generating the Python assertion functions is as follows:
@@ C.1. Evaluating Subsumption
In the prompt above, the assertion_blob represents a string of all assertion functions. Then, the second prompt is as follows:

### v2

@@ 4.3. Results and Discussion
$\spadesuit$ spade Subsumption as a Means for Reducing Redundancy. Several pipelines exhibit a large discrepancy between functions selected in baseline and $\textsc{spade}_{\_}{\text{sub}}$, which occurs when there are many redundant candidates. For example, in the codereviews pipeline’s 8 prompt versions, the developer iterated several times on the instruction to give a clear and concise review, resulting in five assertions that check the same thing (two of which are shown below): $\spadesuit$ spade Since all the five assertions meet the FFR constraint, individually, baseline would select them all, which is undesirable because they all do the same thing, but $\textsc{spade}_{\_}{\text{sub}}$ would select the one most compatible with the FFR constraint, as long as subsumption is assessed correctly. On the flip side, while assessing subsumption, the LLM may not recall all subsumptions, so $\textsc{spade}_{\_}{\text{sub}}$ may have duplicate assertions. For example, the codereviews pipeline contains assertions titled assert_includes_code_improvementv1 and assert_includes_code_improvementv2. $\boldsymbol{\alpha}$ and $\boldsymbol{\tau}$ Threshold Sensitivity. The feasibility of solutions from the ILP solver in spade is dependent on the chosen $\alpha$ and $\tau$ thresholds. If a feasible solution is not found, developers may need to adjust these values in a binary search fashion. In our case, all 9 LLM pipelines yielded feasible solutions with $\alpha=0.6$ and $\tau=0.25$. However, the small size of $E^{\prime}$ makes $\textsc{spade}_{\_}{\text{cov}}$ particularly sensitive to $\alpha$. In the pipelines, we observed that between one and five assertions covered 60% of $E^{\prime}$’s failures. For example, $\textsc{spade}_{\_}{\text{cov}}$ selected only one assertion for the emails pipeline: $\spadesuit$ spade If $E^{\prime}$ is exhaustive of failure modes and representative of the distribution of failures (e.g., for the emails pipeline, most failures are actually due to the response lacking an encouragement to contact the company), $\textsc{spade}_{\_}{\text{cov}}$ might be a satisfactory solution. However, our $E^{\prime}$ datasets clearly were not exhaustive, considering that $\textsc{spade}_{\_}{\text{sub}}$ always chose additional assertions. $\textsc{spade}_{\_}{\text{sub}}$ is less sensitive to $\alpha$, as it explicitly selects assertions based on their potential to cover new failures (i.e., subsumption) without exceeding the FFR, even if the constraint on coverage is no longer tight. FFR Tradeoffs. Considering that the difference between the fraction of functions selected for baseline and $\textsc{spade}_{\_}{\text{sub}}$ is less than 10% for three LLM pipelines, one may wonder if the complexity of $\textsc{spade}_{\_}{\text{sub}}$ is worth it. $\textsc{spade}_{\_}{\text{sub}}$ is generally preferable because baseline fails to consistently meet the FFR threshold $\tau$. We observed that as prompt versions increase, so do the number of assertions, impacting baseline adversely. The worst-case FFR of a set is the sum of individual FFRs, as shown in Equation 2. Hence, with a large number of independent assertions, the total FFR is likely to surpass the threshold. This issue is evident in the fashion and lecturesummaries pipelines, where despite each of the 67 and 32 assertions meeting FFR constraints individually, the total FFR for baseline reaches 88% and 53%, respectively. In practice, if spade were to be deployed in an interactive system, where spade could observe each LLM call in real-time (e.g., as a wrapper around the OpenAI API), the multitude of prompt versions further necessitates filtering assertions based on overall FFR. This underscores the need for the more complex $\textsc{spade}_{\_}{\text{cov}}$ or $\textsc{spade}_{\_}{\text{sub}}$ approaches.
@@ 5. Related Work
Prompt Engineering. For both nontechnical (Zamfirescu-Pereira et al., 2023) and technical users (Si et al., 2022; Parameswaran et al., 2023), prompt engineering is hard for several reasons: small changes in prompt phrasing (Liu et al., 2023; Arora et al., 2022) or the order of instructions or contexts (Lu et al., 2021) can significantly affect outputs. Moreover, as LLMs change under the hood of the API (i.e., prompt drift), outputs can change without developer awareness (Chen et al., 2023). Tools and papers are emerging to aid in prompt management and experimentation, and are even using LLMs to write prompts (Arawjo et al., 2023; Wu et al., 2022b, a; Cheng et al., 2023; Zhou et al., 2022). Moreover, deployed prompts introduce new challenges, like “balancing more context with fewer tokens” and “wrangling prompt output” to meet user-defined criteria (Parnin et al., 2023). Our work doesn’t focus explicitly on helping developers create better prompts, but it could indirectly support developers by helping identify mistakes.
ML and LLM Evaluation. Evaluating and monitoring deployed ML models is known to be challenging (Paleyes et al., 2022; Shankar et al., 2022b). Evaluating LLMs in a deployed setting is even more challenging because LLMs are typically used for generative tasks, where the outputs are free-form (Dou et al., 2021). Some LLM pipeline types, like question-answering with retrieval-augmented generation pipelines (Lewis et al., 2020), can use standardized, automated metrics (Es et al., 2023; Saad-Falcon et al., 2023), but others face challenges due to unknown metrics and lack of labeled datasets (Xiao et al., 2023; Chang et al., 2023; Parnin et al., 2023). Typically, organizations rely on human evaluators for LLM outputs (Gehrmann et al., 2023; Wang et al., 2023c; Parnin et al., 2023), but recent studies suggest LLMs can self-evaluate effectively with detailed “scorecards” (Chan et al., 2023; Zhang et al., 2023; Kim et al., 2023). However, writing these scorecards can be challenging (Parnin et al., 2023), motivating auto-generated evaluators. Recent work (Rebedea et al., 2023; Singhvi et al., 2023; Kim et al., 2023) and industry tools (Guardrails, 2023; Langchain, 2023; Llama Index, 2023) proposes the use of assertions to catch mistakes in LLM pipelines, while requiring the user to select these assertions.
@@
Testing and Validation in ML Pipelines. ML pipelines are hard to manage in production. Much of the literature on ML testing is geared towards validating structured data, through analyzing data quality (Hynes et al., 2017; Breck et al., 2019b; Schelter et al., 2018; Shankar et al., 2023a) or provenance (Schelter et al., 2023; Namaki et al., 2020). Platforms for ML testing typically offer automated experiment tracking and prevention against overfitting (Renggli et al., 2019; Agrawal et al., 2019), as well as tools for data distribution debugging (Grafberger et al., 2022). Model-specific assertions typically require human specification (Kang et al., 2020), or at least large amounts of data to train learned assertions (Kang et al., 2022). LLM chains or pipelines are a new class of ML pipelines, and LLMs themselves can generate assertions with little data. A recent study highlights the difficulty of testing LLM pipelines for “copilot”-like products: developers want to ensure accuracy while avoiding excessive resource use, such as running hundreds of assertions (Parnin et al., 2023)—motivating assertion filtering.
6. Conclusion and Future Work
We introduce a new problem of auto-generating assertions to catch failures in LLM pipelines, as well as spade, our framework for doing so. spade comprises two components: first, it synthesizes candidate assertions, and then it filters them down into a more manageable subset. To synthesize candidate assertions, we analyzed prompt version histories and learned that prompt deltas are often a rich source of requirements and therefore candidate assertions. We developed a taxonomy of prompt deltas for assertion synthesis, demonstrating its value via integration and deployment with LangChain, with over 2000 runs across domains. For the latter problem of candidate assertion filtering, we expressed the selection of an optimal set of assertions, covering most failures while introducing as few false failures as possible as an Integer Linear Program (ILP). We proposed assertion subsumption to cover failures in data-scarce scenarios and incorporated this into our ILP. We also studied the setting where there are no examples and demonstrated that it reduces to a topological sort on the subsumption graph. Our auto-generating assertion system, spade, was evaluated on nine real-world data-generating LLM pipelines. We have made our code and datasets public for further research and analysis.
There are a number of open questions in our effort to make production LLM pipelines more robust. For instance, while meeting developer provided criteria ($\alpha,\tau$) is helpful, sometimes developers would like to examine the generated and selected assertions in a way that helps them make the tradeoffs themselves, motivating a human-in-the-loop interface to assist developers in defining data quality for LLM pipelines. Such an interface could also be a vehicle for getting developers to label examples on the fly. Determining which labeled examples would help best select from the set of assertions is an open question that is reminiscent of active learning. Finally, we could also envision automatically updating assertion sets in deployed pipelines, as new failure modes inevitably arise in production.
@@ Appendix A Notation
$\mathcal{P}$ A prompt template (i.e., a string with placeholders, intended to be submitted to an LLM)
$\mathcal{P}_{\_}i$ The $i$th version of a prompt template
$\Delta\mathcal{P}_{\_}i$ The diff or changes between consecutive versions of a prompt template
$E$ The set of all hypothetical example runs for an LLM pipeline
$E^{\prime}$ A small set of example runs for an LLM pipeline (with labels denoting whether responses were good) that we can observe, where $E^{\prime}\subset E$
$n$ The number of examples in $E^{\prime}$
$e_{\_}i$ The $i$th example in $E^{\prime}$
$f$ An arbitrary assertion function
$F$ The set of all candidate assertions from Section 2
$m$ The number of assertions in $F$
$F^{\prime}$ A subset of $F$ selected as a minimal set of assertions
$G$ Set of functions in $F\setminus F^{\prime}$ not subsumed by $F^{\prime}$ (as defined by Definition 3.3)
$\hat{y}_{\_}i$ Binary indicator of whether example $e_{\_}i$ satisfies all assertions in $F^{\prime}$
$y_{\_}i$ Binary indicator of whether a developer considers the LLM pipeline response for example $e_{\_}i$ good (i.e., success) or bad (i.e., failure)
$\alpha$ Threshold for coverage of failing examples; $F^{\prime}$ should cover at least $\alpha$ of failing examples
$\tau$ Threshold for False Failure Rate (FFR); $F^{\prime}$ should fail no more than $\tau$ good examples
$M$ $n\times m$ matrix representing the results of assertions on examples in $E^{\prime}$
$K$ $m\times m$ matrix representing subsumption relationships between functions
$x_{\_}j$ ILP variable indicating inclusion of $f_{\_}j$ in $F^{\prime}$
$w_{\_}{ij}$ ILP variable representing if $F^{\prime}$ denotes $e_{\_}i$ as a failure
$u_{\_}i$ ILP variable indicating if a failed example is covered
$z_{\_}i$ ILP variable defining if $F^{\prime}$ incorrectly marks a successful example as a failure (false failure)
$r_{\_}j$ ILP variable indicating if a function is subsumed by any function in $F^{\prime}$
$s_{\_}j$ ILP variable representing functions in $F\setminus F^{\prime}$ not subsumed by $F^{\prime}$
@@ Appendix B LangChain Prompt Hub and Taxonomy of Prompt Deltas
$\spadesuit$ spade Let concepts be the parsed categories identified by the previous prompt. The prompt for generating the Python assertion functions is as follows: $\spadesuit$ spade
@@ C.1. Evaluating Subsumption
$\spadesuit$ spade In the prompt above, the assertion_blob represents a string of all assertion functions. Then, the second prompt is as follows: $\spadesuit$ spade